OpenStudy (anonymous):

prove: Csc=1+cot(Ө)/sin(Ө)+cos(Ө)

OpenStudy (anonymous):

\[\csc = 1+ (\cot/(\sin + \cos))? \] or \[\csc = (1 + \cot)/\sin + \cos?\] You need to try and use the website's tools to help us out a bit here, it's impossible to demonstrate without clearly writing it out.

OpenStudy (anonymous):

\[\csc(\theta)=\frac{1+\cot(\theta)}{\sin(\theta)+\cos(\theta)}\]?

OpenStudy (anonymous):

yes exactly..

OpenStudy (anonymous):

ok then this is an easier one. multiply top and bottom by \[\sin(\theta)\]

OpenStudy (anonymous):

\[\csc(\theta)=\frac{1+\cot(\theta)}{\sin(\theta)+\cos(\theta)}\frac{\sin(\theta)}{\sin(\theta)}\] \[\frac{\sin(\theta)+\cos(\theta)}{(\sin(\theta)+\cos(\theta))\sin(\theta)}\] \[=\frac{1}{\sin(\theta)}\] \[=\csc(\theta)\]

OpenStudy (anonymous):

thanks for helping me..